# Optimal Control of an SIR Model with Noncompliance as a Social Contagion

### v1

@@ 3. Optimal Control of our Model
We direct the reader to several classical sources for proofs and discussions regarding the PMP; for example, [2, §5], [29, §4], [42, §1,2], [49, §12]. We note that (3.7) is exactly the original dynamics (2.2). We refer to (3.8) as the costate dynamics, and (3.9) as the optimality condition. The condition $p(T)=0$ in (3.8) is called the transversality condition (discussed in the aforementioned sources), and takes this simplified form because we consider a fixed time, free end-point problem with no terminal cost.
We will use the forward-backward sweep method described in [28, §4] to numerically solve (3.7)-(3.9). We fully detail the algorithm shortly, but first, to make all computations entirely explicit, we write out (3.8) and (3.9) for our specific choice of cost functional (3.2). In our case, the Hamiltonian can be written
(3.10) $\begin{split}\mathcal{H}(x,p,u)&=p_{S}\Big{(}(1-\xi)b-\beta(1-\alpha)S(I+I^{*})-(\overline{\mu}-\mu)SN^{*}+\nu S^{*}-\delta S\Big{)}\\ &\hskip 28.45274pt+p_{I}\Big{(}\beta(1-\alpha)S(I+I^{*})-(\gamma+\eta)I-(\overline{\mu}-\mu)IN^{*}+\nu I^{*}-\delta I\Big{)}\\ &\hskip 28.45274pt+p_{R}\Big{(}(\eta+\gamma)I-(\overline{\mu}-\mu)RN^{*}+\nu R^{*}-\delta R\Big{)}\\ &\hskip 28.45274pt+p_{S^{*}}\Big{(}\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu)SN^{*}-\nu S^{*}-\delta S^{*}\Big{)}\\ &\hskip 28.45274pt+p_{I^{*}}\Big{(}\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-\nu I^{*}-\delta I^{*}\Big{)}\\ &\hskip 28.45274pt+p_{R^{*}}\Big{(}\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-\nu R^{*}-\delta R^{*}\Big{)}\\ &\hskip 28.45274pt+c_{1}(I+I^{*})+c_{2}N^{*}+c_{3}\frac{\alpha^{2}}{2}+c_{4}\frac{\eta^{2}}{2}+c_{5}\frac{\mu^{2}}{2}+c_{6}\frac{\nu^{2}}{2}.\end{split}$
Thus, (3.8) reads
(3.11) $\begin{split}\frac{dp_{S}}{dt}&=\beta(1-\alpha)(I+I^{*})(p_{S}-p_{I})+(\overline{\mu}-\mu)N^{*}(p_{S}-p_{S^{*}})+\delta p_{S},\\ \frac{dp_{I}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+(\gamma+\eta)(p_{I}-p_{R})+\delta p_{I}-c_{1}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)N^{*}(p_{I}-p_{I^{*}}),\\ \frac{dp_{R}}{dt}&=(\overline{\mu}-\mu)N^{*}(p_{R}-p_{R^{*}})+\delta p_{R},\\ \frac{dp_{S^{*}}}{dt}&=\beta(I+I^{*})(p_{S^{*}}-p_{I^{*}})+\nu(p_{S^{*}}-p_{S})+\delta p_{S^{*}}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)},\\ \frac{dp_{I^{*}}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+\gamma(p_{I^{*}}-p_{R^{*}})+\nu(p_{I^{*}}-p_{I})-c_{1}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{I^{*}},\\ \frac{dp_{R^{*}}}{dt}&=\nu(p_{R^{*}}-p_{R})+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{R^{*}}-c_{2},\end{split}$
along with zero terminal condition $p_{X}(T)=0$ for each $X\in\{S,I,R,S^{*},I^{*},R^{*}\}.$
Because the Hamiltonian is quadratic in the control variables $u=(\alpha,\eta,\mu,\nu)$, one can easily resolve the optimality condition (3.9) by simply minimizing a quadratic function and projecting back onto the admissible set of control values. Suppressing dependence on $t$, once the state and costate variables are known, optimal controls are given by
(3.12) $\begin{split}\alpha&=\min\left\{\max\left\{\frac{\beta S(I+I^{*})(p_{I}-p_{S})}{c_{3}},\,0\right\},\,1\right\},\\ \eta&=\min\left\{\max\left\{\frac{I(p_{I}-p_{R})}{c_{4}},\,0\right\},\,\overline{\eta}\right\},\\ \mu&=\min\left\{\max\left\{\frac{[S(p_{S^{*}}-p_{S})+I(p_{I^{*}}-p_{I})+R(p_{R^{*}}-p_{R})]N^{*}}{c_{5}},\,0\right\},\,\overline{\mu}\right\},\\ \nu&=\min\left\{\max\left\{\frac{S^{*}(p_{S^{*}}-p_{S})+I^{*}(p_{I^{*}}-p_{I})+R^{*}(p_{R^{*}}-p_{R})}{c_{6}},\,0\right\},\,\overline{\nu}\right\}.\end{split}$
Given this, the forward-backward sweep method is detailed in algorithm 1. We include some basic implementation notes. For the initial guess, we always use $u^{(0)}\equiv 0$. Step (iv) in the iteration is a relaxation step which may not be necessary but, empirically, accelerates convergence. For convergence criterion, we halt the iteration when the relative change in control variables (measured in supremum norm) is below some specified tolerance. Using the suggestion of [28, §4], we rearrange the condition slightly so as to avoid division by zero in the case that zero control is optimal. Thus our convergence criterion is
(3.13) $(\text{tol})\cdot\|\alpha^{(k+1)}\|_{\infty}-\|\alpha^{(k+1)}-\alpha^{(k)}\|_{\infty}\geq 0,$
where “tol” is the convergence tolerance. While (3.13) is written for $\alpha$, we enforce this criterion for each of $\alpha,\eta,\mu,\nu$ and only halt the iteration once all of them satisfy the condition.
Algorithm 1 Forward-Backward Sweep Method for Resolving Optimal Controls
Input: initial condition $x_{0}=(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})\in{\mathbb{R}}^{6}_{+}$, parameter values $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$, cost weights $c_{1},\ldots,c_{6}$, a relaxation parameter $\kappa\in[0,1]$ and an initial guess for the control map $u^{(0)}=(\alpha^{(0)},\eta^{(0)},\mu^{(0)},\nu^{(0)})$. Set $k=0$ and do the following:
@@
  (i) Resolve $x^{(k+1)}=(S^{(k+1)},I^{(k+1)},R^{(k+1)},(S^{*})^{(k+1)},(I^{*})^{(k+1)},(R^{*})^{(k+1)})$ as the solution of (2.2) with initial condition $x_{0}$, using control variables $u^{(k)}$
  (ii) Resolve $p^{(k+1)}=(p_{S}^{(k+1)},p_{I}^{(k+1)},p_{R}^{(k+1)},p_{S^{*}}^{(k+1)},p_{I^{*}}^{(k+1)},p_{R^{*}}^{(k+1)})$ as the solution of (3.11) with zero terminal condition, using control variables $u^{(k)}$ and state variables $x^{(k+1)}$
  (iii) Define $u^{*}=(\alpha^{(k+1)},\eta^{(k+1)},\mu^{(k+1)},\nu^{(k+1)})$ from (3.12) using state variables $x^{(k+1)}$ and costate variables $p^{(k+1)}$
  (iv) Define $u^{(k+1)}=(1-\kappa)u^{(k)}+\kappa u^{*}$ and increment $k$
until convergence criteria are met
return $u=u^{(k)}$, $x=x^{(k)}$, $p=p^{(k)}$: respectively the optimal control, state, and costate trajectories
@@ 4. Results & Discussion
We implement algorithm 1 in MATLAB and demonstrate the results in several different cases. Specifically, we are interested in observing the differing behavior of optimal control maps in response to changes in the policy-maker’s preferences, expressed via the weights $c_{1},\ldots,c_{6}$ from (3.1). For all simulations, we use relaxation parameter $\kappa=0.8$ and a time horizon of $T=100$, and for the sake of numerically resolving the solutions of (2.2) and (3.11), we use an explicit Euler scheme with uniform discretization of $[0,T]$ with step size $\Delta t=0.1$, and a convergence tolerance of $\text{tol}=10^{-3}.$ The code which generated all of these results is available online. [^1] https://github.com/chparkinson/Opt_Control_SIR_With_Noncompliance_As_Social_Contagion
@@
What remains is to specify initial conditions and the parameters $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$. To enhance readability, we include a description of each of these parameters and the values they take in each scenario. In all scenarios, we take $b=\delta$, and initial condition $(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})=(0.69,0.01,0,0.29,0.01,0)$, indicating that at the outset of the epidemic, 2% of the population is infected, and 30% of the population is noncompliant. Because $b=\delta$ and the initial condition has total population $1$, by (2.6), the total population will be $1$ for all time, so individual compartments can be interpreted as percentages of the total population.
@@
Before displaying results, we briefly discuss the bearing of the controls. Recall that the cost (3.2) for the policy-maker is split into three parts: (C1) cost associated with infections (weighted by the constant $c_{1}$), (C2) costs associated with noncompliance (weighted by the constant $c_{2}$) and (C3) cost associated with controls (weighted by constants $c_{3},\ldots,c_{6}$). By proposition 2.2, any use of controls will help decrease cost (C1) by slowing the spread of infections. This is the only effect that $\alpha$ and $\eta$ can have. The controls $\mu$ and $\nu$ obliquely lower infection rate and thus reduce cost (C1), but more explicitly decrease cost (C2) by decreasing the noncompliant population. Of course, any use of control increases cost (C3), meaning that the policy-maker has a trade-off to consider. We mention specifically that the absolute size of the weights is not important, but rather their size relative to each other. Accordingly, in much of the below we fix the control cost weights $c_{3},\ldots,c_{6}=1$ and vary the cost weights $c_{1},c_{2}$ associated with total infections and total amount of noncompliance, respectively. In In certain cases, we would like to “turn off” individual controls; to do so, one can simply set that control to zero (this could equivalently be accomplished by sending the control’s cost weight to infinity). Lastly, as baseline values for infection and noncompliance cost weights, we choose $c_{1}=1$ and $c_{2}=0.1$. This causes the cost due to infection and the cost due to noncompliance to be roughly equal in the uncontrolled case. The reason that $c_{2}$ must be significantly smaller is that with no controls, noncompliance is a permanent state, whereas infection is not, which means that noncompliance has the potential to contribute much more to the total cost.
@@
$\overline{\mu}$ baseline infection rate for noncompliance 0.1 0.2 0.5
@@
$c_{2}$ cost weight on infections var. var. 0.1
$c_{3},\ldots,c_{6}$ cost weights on controls 1 1 1
Figure 1. Table of parameters for each scenario. Recall in particular that $\overline{\eta},\overline{\mu},\overline{\nu}$ are upper bounds for the control variables $\eta,\mu,\nu$, which represent increase in recovery rate due to seeking treatment, decrease in spread rate of noncompliance due to public health initiatives, and recovery rate for noncompliance due to educational campaigns, respectively. Not mentioned is the control variable $\alpha$, representing decrease in infection rate among the compliant susceptible population due to nonpharmaceutical intervention measures, since $\alpha$ will take values in $[0,1].$

### v2

@@ 3. Optimal Control of our Model
We direct the reader to several classical sources for proofs and discussions regarding the PMP; for example, [3, §5], [34, §4], [48, §1,2], [58, §12]. We note that (3.10) is exactly the original dynamics (2.2). We refer to (3.11) as the costate dynamics, and (3.12) as the optimality condition. The condition $p(T)=0$ in (3.11) is called the transversality condition (discussed in the aforementioned sources), and takes this simplified form because we consider a fixed time, free end-point problem with no terminal cost.
In the ensuing section, we describe numerically resolving optimal controls via the sequential quadratic Hamiltonian method [10].
4. Numerical Methods
In this section, we consider the numerical resolution of optimal control maps via the sequential quadratic Hamiltonian (SQH) method [10]. Because this is a fairly new numerical method, and to make this manuscript as self-contained as possible, we discuss some basic facets of the method and prove convergence of the iteratively defined control maps. A full discussion of the method with worked examples can be found in [10].
The SQH method is an iterative method for resolving control maps in a variety of scenarios involving optimal control of ordinary or partial differential equations. The basic strategy is to successively update the approximation of the control maps using quadratic perturbations to the true Hamiltonian (3.9). Specifically, for $\varepsilon>0$, defined the auxiliary Hamiltonian
(4.1) $\mathcal{H}_{\varepsilon}(x,p,u,\overline{u})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{H}(x,p,\overline{u})+\varepsilon\left|u-\overline{u}\right|^{2}_{2}.$
The quadratic term above introduces relaxation, so that solving
$u^{\text{new}}(t)=\operatorname*{argmin}_{w\in U_{\text{ad}}}\mathcal{H}_{\varepsilon}(x(t),p(t),u(t),w)$
results in $u^{\text{new}}$ which balances minimizing the Hamiltonian with a penalty for straying too far from the currently employed control strategy $u$.
Given this definition, the full SQH algorithm is detailed in algorithm 1. The basic strategy of the algorithm is to successively update $x,p$ and $u$ by solving the state and co-state equations and resolving controls via (4.1). One significant novelty is the inclusion of step (v) in the iteration, whereupon one checks the condition
(4.2) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau.$
If this condition is met, it is determined that the new controls $u^{(k+1)}$ have achieved a sufficient decrease in the cost functional. In this case, the new controls $u^{(k+1)}$ are accepted, $\varepsilon$ is decreased by setting $\varepsilon\leftarrow\zeta\varepsilon$ for some $\zeta\in(0,1)$, and the iteration continues with the updated auxiliary Hamiltonian. If the condition is not met, there was not sufficient decrease in the cost functional, so $\varepsilon$ is increased by setting $\varepsilon\leftarrow\lambda\varepsilon$ for some $\lambda>1$ and $u^{(k+1)}$ is recomputed before continuing. Here $\lambda,\zeta$ are tunable parameters, as is $\rho$ in (4.2), which determines the level to which $u^{(k+1)}$ must decrease the cost functional to be considered acceptable.
Algorithm 1 Sequential Quadratic Hamiltonian Method for Resolving Optimal Control Maps
(1) Input: initial approximation $u^{0}$, maximum number of iterations $k_{max}$, tolerance $\kappa>0$, and hyperparamaters $\varepsilon>0$, $\lambda>1$, $\rho>0$, and $\zeta\in\left(0,1\right)$.
(2) Set $k=0$ and compute the solution $x^{(0)}=(S^{(0)},I^{(0)},R^{(0)},S^{*,(0)},I^{*,(0)},R^{*,(0)})$ to (2.2) with control $u=u^{(0)}$.
(3) Perform the following iteration.
@@ 4. Numerical Methods
  (i) Compute the costate dynamics $p^{(k)}$ obeying (3.11) with $x=x^{(k)}$ and $u=u^{(k)}$.
  (ii) Compute $u^{(k+1)}$ via:
$u^{(k+1)}(t)=\operatorname*{argmin}_{w\in U_{\text{ad}}}\,\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),w\right),\,\,\,\,\,\,\,t\in[0,T].$
  (iii) Compute the solution $x^{(k+1)}$ to the (2.2) with control $u=u^{(k+1)}.$
  (iv) Compute $\tau=\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$.
  (v) if: $\left\{J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau\right\}$ then set $\varepsilon\leftarrow\zeta\varepsilon$ and continue to step (vi),
     else if: $\left\{J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)>-\rho\,\tau\right\}$ then set $\varepsilon\leftarrow\lambda\varepsilon$ and go to step (ii).
  (vi) Set $k\leftarrow k+1.$
until $k=k_{\text{max}}$ or $\tau<\kappa$
return $(x^{*},u^{*})$ - the most recently computed state and control trajectories
We prove a basic convergence result for algorithm 1, after establishing a lemma.
Lemma 4.1.
Assume that the state cost function $r$ from (3.1) has a Lipschitz continuous gradient. Let $u,v\in\mathscr{U}_{\text{ad}}$ be two control maps with associated solutions $x=\mathcal{S}(u)$ and $y=\mathcal{S}(v)$ of (2.2). Then there is a constant $c>0$ such that
(4.3) $J(x,u)-J(y,v)\leq\int^{T}_{0}\Big(\mathcal{H}(x(t),p(t),u(t))-\mathcal{H}(x(t),p(t),v(t))\Big)dt+c\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
where $p$ is the co-state trajectory corresponding to $(x,u)$.
Proof.
We compute, dropping depending on $t$ for brevity,
(4.4) $J(x,u)-J(y,v)$ $=\int^{T}_{0}\Big(r(x)+\ell(u)-r(y)-\ell(v)\Big)dt$
$=\int^{T}_{0}\Big([\mathcal{H}(x,p,u)-\mathcal{H}(y,p,v)]+\left\langle p,f(y,v)-f(x,u)\right\rangle\Big)dt$
$=\int^{T}_{0}[\mathcal{H}(x,p,u)-\mathcal{H}(x,p,v)]dt$
$\hskip 56.9055pt+\int^{T}_{0}\Big([\mathcal{H}(x,p,v)-\mathcal{H}(y,p,v)]+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\Big)dt.$
In the last equality of (4.4), the first term on the right hand side is as it appears in (4.3), so we work on the last line, which we call $I$. We see
(4.5) $I$ $=\int^{T}_{0}\Big([\mathcal{H}(x,p,v)-\mathcal{H}(y,p,v)]+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\Big)dt$
$=\int^{T}_{0}\left(\int^{1}_{0}-\left\langle\nabla_{x}\mathcal{H}(x+\lambda(y-x),p,v),y-x\right\rangle d\lambda+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\right)dt.$
Now integrating by parts to pass the derivative onto $p$, the boundary terms vanish since $x(0)=y(0)$ and $p(T)=0$. So using (3.11), (4.5) becomes
(4.6) $I=\int^{T}_{0}\int^{1}_{0}\left\langle\nabla_{x}\mathcal{H}(x,p,u)-\nabla_{x}H(x+\lambda(y-x),p,v),y-x\right\rangle d\lambda dt.$
Now $\nabla_{x}\mathcal{H}=(J_{f}(x,u))^{T}p+\nabla r$ where $J_{f}$ is the Jacobian matrix of $f$. Since $f$ is smooth (and its arguments remain bounded) and since $\nabla r$ is Lipschitz by assumption, $\nabla_{x}\mathcal{H}$ is Lipschitz. Thus (4.6) implies
(4.7) $I$ $\leq L\int^{T}_{0}\int^{1}_{0}(\lambda\left|x-y\right|^{2}_{2}+\left|u-v\right|_{2}\left|x-y\right|_{2})d\lambda dt$
$\leq\frac{L}{2}\|x-y\|^{2}_{(L^{2}[0,T])^{6}}+L\|x-y\|_{(L^{2}[0,T])^{6}}\|u-v\|_{(L^{2}[0,T])^{4}}.$
Finally, lemma 3.2 tells us that the control-to-state map is Lipschitz, whereupon (4.7) yields
$I\leq c\|u-v\|^{2}_{(L^{2}[0,T])^{2}}$
and inserting this last bound into (4.4) concludes the proof. ∎
Theorem 4.2.
Under the assumptions of lemma 4.1, let $\left(x^{(k)},u^{(k)}\right)$ and $\left(x^{(k+1)},u^{(k+1)}\right)$ be successive iterates generated by the SQH method (algorithm 1). Then there is $c>0$ such that
(4.8) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-(\varepsilon-c)\,\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}.$
for the current value of $\varepsilon>0$ in algorithm 1. In particular, this implies
(4.9) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau$
for $\varepsilon\geq\rho+c$ and $\tau=\|u^{k+1}-u^{k}\|^{2}_{(L^{2}[0,T])^{3}}$.
Proof.
In step (ii), we resolve $u^{(k+1)}$ such that
$\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),u^{(k+1)}(t)\right)\leq\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),w\right),$
for all $w\in U_{\text{ad}}\text{ and }t\in[0,T].$ In particular, choosing $w=u^{(k)}(t)$ at each $t\in[0,T]$ (and then dropping the argument $t$ for brevity), we see
$\mathcal{H}_{\varepsilon}\left(x^{(k)},p^{(k)},u^{(k)},u^{(k+1)}\right)\leq\mathcal{H}_{\varepsilon}\left(x^{(k)},p^{(k)},u^{(k)},u^{(k)}\right)=\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big).$
Inserting the definition of $\mathcal{H}_{\varepsilon}$, this yields
$\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k+1)}\Big)-\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big)\leq-\varepsilon\big|u^{(k+1)}-u^{(k)}\big|^{2}_{2}.$
Then from lemma 4.1, we have
$J\Big(x^{(k+1)},u^{(k+1)}\Big)-J\Big(x^{(k)},u^{(k)}\Big)$ $\leq\int^{T}_{0}\Big(H\Big(x^{(k)},p^{(k)},u^{(k+1)}\Big)-\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big)\Big)dt$
$\hskip 85.35826pt+c\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$
$\leq-(\varepsilon-c)\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$
as desired. ∎
Recall, if sufficient decrease of the cost functional is not achieved, $\varepsilon$ is successively increased. Theorem 4.2 tells us that once $\varepsilon\geq\rho+c$, we are guaranteed to achieve sufficient decrease of the cost functional. Given this, we prove that the numerical iterates do indeed converge.
Theorem 4.3.
The sequence of iterates $\{(x^{(k)},u^{(k)})\}$ created by the SQH algorithm satisfy the following.
  • (a)
    $\{J\Big(x^{(k)},u^{(k)}\Big)\}$ decreases in $k$, and thus converges to some $J^{*}\geq\min_{u\in\mathscr{U}_{\text{ad}}}J(x,u)$
  • (b)
    $\|u^{(k+1)}-u^{(k)}\|_{(L^{2}[0,T])^{4}},\,\|x^{(k+1)}-x^{(k)}\|_{(L^{2}[0,T])^{6}}\to 0$ as $k\to\infty$
Proof.
Monotonicity of $J(x^{(k)},u^{(k)})$ follows directly from (4.9) following the observation above that eventually $\varepsilon$ will be increased enough that $\varepsilon\geq\rho+c$. Convergence then follows by the monotone convergence theorem, so (a) is proven.
For (b), we rearrange (4.9) to see that when $u^{(k+1)}$ is resolved,
$\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}\leq\frac{1}{\rho}\Big(J\Big(x^{(k)},u^{(k)}\Big)-J\Big(x^{(k+1)},u^{(k+1)}\Big)\Big).$
Thus we have the telescoping sum
$\sum^{K}_{k=0}\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}\leq\frac{1}{\rho}\Big(J\Big(x^{(0)},u^{(0)}\Big)-J\Big(x^{(K+1)},u^{(K+1)}\Big)\Big).$
As $K\to\infty$ the right hand side above converges, so this shows that the left hand side is a convergent series of positive terms, implying that the summands go to zero. That $\|x^{(k+1)}-x^{(k)}\|_{(L^{2}[0,T])^{6}}\to 0$ as well follows by the Lipschitz continuity of the control-to-state map proven in lemma 3.2. ∎
With this, we move on to simulations and discussion wherein we define a cost functional discuss the optimal control strategies that result in the case of different policy-maker preferences.
@@ 5. Results & Discussion
We implement algorithm 1 in MATLAB and demonstrate the results in several different cases. Specifically, we consider the cost functional (3.1) with
(5.1) $r(x)$ $=c_{1}(I+I^{*})+c_{2}(S^{*}+I^{*}+R^{*}),$
$\ell(u)$ $=d_{1}(\left|\alpha\right|+\left|\eta\right|+\left|\mu\right|+\left|\nu\right|)+\frac{d_{2}}{2}(\alpha^{2}+\eta^{2}+\mu^{2}+\nu^{2}).$
where $c_{1},c_{2},d_{1},d_{2}$ are nonnegative weights that express the policy-maker’s preferences. We are interested in observing the differing behavior of optimal control maps in response to changes in the policy-maker’s preferences. That is, in each scenario below, we fix these weights and use the SQH method to resolve approximate optimal control maps.
For all simulations we set $T=100$, and for the sake of numerically resolving the solutions of (2.2) and (3.11), we use an explicit Euler scheme with uniform discretization of $[0,T]$ with step size $\Delta t=0.1$.
@@
What remains is to specify initial conditions and the parameters $b,\delta,\beta,\gamma,\overline{\eta},\xi,\mu^{*},\overline{\mu},\overline{\nu}$. To enhance readability, we include a description of each of these parameters and the values they take in each scenario. In all scenarios, we take $b=\delta$, and initial condition $(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})=(0.69,0.01,0,0.29,0.01,0)$, indicating that at the outset of the epidemic, 2% of the population is infected, and 30% of the population is noncompliant. Because $b=\delta$ and the initial condition has total population $1$, by (2.3), the total population will be $1$ for all time, so individual compartments can be interpreted as percentages of the total population.
@@
Before displaying results, we briefly discuss the bearing of the controls. Recall that the cost (3.1) for the policy-maker is split into three parts: (C1) cost associated with infections (weighted by the constant $c_{1}$), (C2) costs associated with noncompliance (weighted by the constant $c_{2}$) and (C3) cost associated with controls (weighted by constants $d_{1},d_{2}$). By proposition 2.2, any use of controls will help decrease cost (C1) by slowing the spread of infections. This is the only effect that $\alpha$ and $\eta$ can have. The controls $\mu$ and $\nu$ obliquely lower infection rate and thus reduce cost (C1), but more explicitly decrease cost (C2) by decreasing the noncompliant population. Of course, any use of control increases cost (C3), meaning that the policy-maker has a trade-off to consider. We mention specifically that the absolute size of the weights is not important, but rather their size relative to each other. Accordingly, in much of the below we fix the control cost weights $d_{1}=d_{2}=0.1$ and vary the cost weights $c_{1},c_{2}$ associated with total infections and total amount of noncompliance, respectively. In certain cases, we would like to “turn off” individual controls; to do so, one can simply set the upper bound for that control to zero (this could equivalently be accomplished by sending the control’s cost weight to infinity). Lastly, as baseline values for infection and noncompliance cost weights, we choose $c_{1}=1$ and $c_{2}=0.1$. This causes the cost due to infection and the cost due to noncompliance to be roughly equal in the uncontrolled case. The reason that $c_{2}$ must be significantly smaller is that with no controls, noncompliance is a permanent state, whereas infection is not, which means that noncompliance has the potential to contribute much more to the total cost. Finally, we emphasize that for each separate plot of controls below, we are changing certain and resolving optimal controls again using the SQH method (i.e., none of these plots contain synthetically designed controls; all of the controls presents are solutions of an optimal control problem with different parameters and/or cost weights).
@@
$\overline{\alpha}$ max. achievable decrease in mixing due to compliance 0.5 0.5 0.5
@@
$\mu^{*}$ baseline infection rate for noncompliance 0.1 0.2 0.3
$\overline{\mu}$ max. achievable decrease in inf. rate of noncomp. 0.05 0.1 0.15
@@
$c_{2}$ cost weight on noncompliance var. 0.01 0.05
$d_{1}$ $L^{1}$ cost weights on controls 0.1 0.1 0.1
$d_{2}$ $L^{2}$ cost weight on controls var. 0.1 0.1
Figure 1. Table of parameters for each scenario. Recall in particular that $\overline{\eta},\mu^{*},\overline{\nu}$ are upper bounds for the control variables $\eta,\mu,\nu$, which represent increase in recovery rate due to seeking treatment, decrease in spread rate of noncompliance due to public health initiatives, and recovery rate for noncompliance due to educational campaigns, respectively. Not mentioned is the control variable $\alpha$, representing decrease in infection rate among the compliant susceptible population due to nonpharmaceutical intervention measures, since $\alpha$ will take values in $[0,1].$